OpenStudy (anonymous):

the frosty ice/cream shop sells sundaes for $2 and banana splits for $3. On a hot summer day, the shop sold 8 more sundaes than banana splits and made $156.

OpenStudy (anonymous):

The question is incomplete.. Complete it first.. What we have to find??

OpenStudy (anonymous):

@skelley1998 , just tell me what we have to find here, you are just written two statements, where is that Interrogative Statement I am looking for??

OpenStudy (anonymous):

See, cost of one sundae is $2 and cost of one banana split is $3.. Let us assume that the shop on hot summer day sold 'x' banana split and 'y' sundaes.. Okay??

OpenStudy (anonymous):

If number of banana split sold is x, shop sold 8 sundae more than banana split that is (x+8). So, total cost of selling sundaes = 2*(x+8) and total cost of selling banana split = 3*x And you are given with total cost of selling as 156. So, \[2(x+8) + 3x = 156\] Can you solve it for x now??